OpenStudy (anonymous):

Find an equation for the tangent to the curve at the given point. y=x^2-x,(4,12) Any ideas on how to solve this problem?

OpenStudy (anonymous):

take the derivative first

OpenStudy (anonymous):

2x?

OpenStudy (anonymous):

y'=2x-1

OpenStudy (anonymous):

so at that point y'(4) will give you the slope of the line

OpenStudy (abb0t):

@Brandon77 is correct. Find the slope and then plug in the given point for x. That's your slope (m) at that given point. Now that you have your slope. Put it into the slope equation.

OpenStudy (anonymous):

Ah, okay. Then I use y = mx +b and find b.

OpenStudy (anonymous):

you got it

OpenStudy (abb0t):

I would leave it as is unless they ask you to re-write it in standard form.. I mean, i'm assuming this is a Calculus course and it's assumed that you've mastered your algebra.

OpenStudy (anonymous):

You can also put it into \[y-y_{1}=m(x-x _{1}) \] and you wouldn't haven't to find b.

OpenStudy (anonymous):

Okay. Either way the final answer is y=7x-16.

OpenStudy (anonymous):

You are correct.